Heather is collecting dimes she saves one dime on the first day two dimes on the second day and three dimes on the third day. If this pattern continues how much money will heather have saved at the end of 30 days?
a) $45.50
b)$46.00
c)$46.10
d)$46.50
e)$47.50


Would you mind to show me how you got it? Thank you

Answer by jim_thompson5910(35256)

You have this arithmetic sequence
1, 2, 3, 4, ..., 28, 29, 30

The first term is 1, the last term is 30. There are 30 terms. So,

a1 = 1
an = 30
n = 30

Use the arithmetic sequence partial sum formula

Sn = n*(a1+an)/2

to get

Sn = n*(a1+an)/2

Sn = n*(1+30)/2

S30 = 30*(1+30)/2

S30 = 465

So she has 465 dimes. Convert to dollars

(465 dimes)*(1 dollar/10 dimes) = 465/10 = 46.50 dollars

So the answer is choice D.
------------------------------------------------------------------------------------------------------------------------

If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).

Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.

Jim
------------------------------------------------------------------------------------------------------------------------